\begin{thebibliography}{BMT}
\bibitem{BV} Bardelli, F.; Verdi, L. Sulle variet\`a razionali di codimensione 2 e di grado 3 in ${\bf P}^n({\bf C})$ come intersezioni complete insiemistiche. Centro di Analisi Globale, Firenze (1979).
\bibitem{B0} Barile, M. Arithmetical ranks of ideals associated to symmetric and alternating matrices. J.~Algebra 176:59--82. 
\bibitem{B1} Barile, M. On the Number of Equations Defining Certain Varieties. Manuscripta Math. 91:483--494.
\bibitem{B2} Barile, M. On the computation of arithmetical ranks. Int.~J.~Pure Appl.~Math.~17:143--161.
\bibitem{BL} Barile, M.; Lyubeznik, G. Set-theoretic complete intersections in characteristic p. Proc.~Amer.~Math.~Soc.~133:3199--3209.
\bibitem{BM1} Barile, M.; Morales, M. On Certain Algebras of Reduction Number One. J.~Algebra 206:113--128. 
\bibitem{BM2} Barile, M.; Morales, M. On the Equations Defining Minimal Varieties. Comm.~Algebra 28:1223--1239. 
\bibitem{BM3} Barile, M.; Morales, M. On Unions of Scrolls Along Linear Spaces. Rend.~Sem.~Mat.~Univ.~Padova 111:161--178. 
\bibitem{BMT} Barile, M.; Morales, M.; Thoma, A. Set-theoretic complete intersections on binomials. Proc.~Amer.~Math.~Soc.~130:1893--1903.
\bibitem{Br} Bresinsky, H. Monomial Gorenstein curves in ${\bf A}^4$ are set-theoretic complete intersections. Manuscripta Math.~27:353--358.
\bibitem{BS} Bruns, W.; Schw\"anzl, R. The number of
equations defining a determinantal variety. Bull.~London Math.~Soc.~22:439--445.
\bibitem{cocoa} CoCoATeam,
  CoCoA, a system for doing Computations in Commutative Algebra.
  Available at \underline{http://cocoa.dima.unige.it}.
\bibitem{EG} Eisenbud, D., Goto, S. Linear Free Resolutions and Minimal Multiplicity. J.~Algebra 88:89--133. 
\bibitem{El} Eliahou, S. (1984) Id\'eaux de d\'efinition des courbes monomiales. In: Greco, S.; Strano, R. ed. {\it Complete Intersections}. Lecture Notes in Mathematics \textbf{1092} {\it Springer}, pp. 229--240.
\bibitem{E} Eto, K. Almost complete intersection monomial curves in ${\bf A}^4$. Comm.~Algebra 22:5325--5342.
\bibitem{G} Gim\'enez, Ph. \'Etude de la fibre sp\'eciale de l'\'eclatement d'une vari\'et\'e monomiale en codimension deux. Dissertation. Institut Fourier, Universit\'e de Grenoble I, 1993. 
\bibitem{GMS} Gim\'enez, Ph.; Morales, M.; Simis, A., L'analytic spread de l'id\'eal de d\'efinition d'une vari\'et\'e monomiale de codimension deux est toujours inf\'erieur ou \'egal \`a trois. (The analytic spread of the ideal of definition of a codimension 2 monomial variety is always less or equal to 3). C.~R.~Acad.~Sci., Paris, S\'er.~I 319:703--706.
\bibitem{H} Harris, J. (1992) {\it Algebraic Geometry.} Springer. 
\bibitem{Hu} Huneke, C. (2004) Lectures on local cohomology (with an appendix by Amelia Taylor). Available at \underline{http://www.math.ku.edu/\~{}huneke/Vita/chicago-lc.pdf} 
\bibitem{K} Katsabekis, A. Projection of cones and the arithmetical rank of toric varieties. Preprint. \underline{arXiv:math.AC/0410264} 
\bibitem{L} Lyubeznik, G. (1984) On the local cohomology modules $H^i_{\cal A}(R)$ for ideals ${\cal A}$ generated by monomials in an $R$-sequence. In: Greco, S.; Strano, R., ed. {\it Complete Intersections}. Lect.~Notes Math.~{\bf 1092}. {\it Springer}, pp. 214--220.
\bibitem{Mi} Milne, J.S. Lectures on \'etale cohomology. Available at \underline{http://www.jmilne.org}.
\bibitem{N} Newstead, P.E. Some subvarieties of Grassmannians of codimension 3. Bull.~London Math.~Soc.~12:176--182.
\bibitem{P} Patil, D.P. Certain monomial curves are set-theoretic complete intersections. Manuscripta Math.~68:399--404. 
\bibitem{PS} Peskine, C.; Szpiro, L. Dimension finie et cohomologie locale. Publ.~Math.~IHES 42:47--119. 
\bibitem{RV1} Robbiano, L.; Valla, G. (1983) Some curves in $P^3$ are set-theoretic complete intersections. In: Ciliberto, C.; Ghione, F.; Orecchia, F., ed. {\it Algebraic Geometry-Open problems}.
Lect.~ Notes Math.~ {\bf 997}. {\it Springer}, pp. 391--399. 
\bibitem{RV2} Robbiano, L.; Valla, G. On set-theoretic complete intersections in the projective space. Rend.~Semin.~Mat.~Fis.~Milano LIII:333--346.
\bibitem{SV} Schmitt, Th.; Vogel, W., Note on Set-Theoretic Intersections of Subvarieties of Projective Space. Math.~Ann.~245:247--253.
\bibitem{X} Xamb\'o, S. On projective varieties of minimal degree.  Collect.~Math.~32:149--163.
\bibitem{Z} Yan, Z. An \'etale analog of the Goresky-MacPherson formula for subspace arrangements. J.~Pure Appl.~Algebra 146:305--318. 

\end{thebibliography}

\documentclass{article}
\catcode`\@=11
\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\newtheorem{example}{Example}
\newtheorem{remark}{Remark}

\def\demo{\noindent{\bf Proof .-}}
\def\section{\@startsection {section}{1}{\z@}{-3.5ex plus -1ex
minus-.2ex}{2.3ex plus .2ex}{\normalsize\bf}}

\def\bz{\hbox{\it Z\hskip -4pt Z}}

\newcommand{\het}{H_{\rm et}}
\newcommand{\hc}{H_{\rm c}}

\begin{document}
\begin{center}
{\Large\bf \textsc{Certain minimal varieties are set-theoretic complete intersections}}\footnote{MSC 2000: 14M10, 14M20}
\end{center}
\vskip.5truecm
\begin{center}
{Margherita Barile\footnote{Partially supported by the Italian Ministry of Education, University and Research.}\\ Dipartimento di Matematica, Universit\`{a} di Bari,Via E. Orabona 4,\\70125 Bari, Italy}
\end{center}
\vskip1truecm
\noindent
{\bf Abstract} We present a class of  homogeneous ideals which are generated by monomials and binomials of degree two and are set-theoretic complete intersections. This class includes certain reducible varieties of minimal degree and, in particular,  the presentation ideals of the fiber cone algebras of monomial varieties of codimension two. 
\vskip0.5truecm
\noindent
Keywords: Minimal variety, rational normal scroll, set-theoretic complete intersection, fiber cone. 

\section*{Introduction and Preliminaries}
Let $K$ be an algebraically closed field, and let $R$ be a polynomial ring  in $N$ indeterminates over $K$. Let $I$ be a proper reduced ideal of $R$ and consider the variety $V(I)$ defined in the affine space $K^N$ (or in the projective space ${\bf P}_K^{N-1}$, if $I$ is homogeneous and different from the maximal irrelevant ideal) by the vanishing of all polynomials in $I$. By Hilbert Basissatz there are  finitely many polynomials $F_1,\dots, F_s\in R$ such that $V(I)$ is defined by the equations $F_1=\cdots= F_s=0$. By Hilbert Nullstellensatz this is equivalent to the ideal-theoretic condition

\begin{equation}\label{1}I=\sqrt{(F_1,\dots, F_s)}.\end{equation}
\noindent
Suppose $s$ is minimal with respect to this property. It is well known that height\,$I\leq s$. If equality holds, $I$ is called a {\it set-theoretic complete intersection} (s.t.c.i.) on $F_1,\dots, F_s$. \newline
Exhibiting significant examples of s.t.c.i. ideals (or, more generally, determining the minimum number of equations defining given varieties, the so-called {\it arithmetical rank}) is one of the most difficult problems in algebraic geometry. The main problem is finding polynomials $F_1,\dots, F_s$ fulfilling (\ref{1}), which need not be part of a minimal generating system for 
$I$. This task cannot be accomplished by a general constructive method. There are, however, a few results that allow us to settle several special cases. One of these is due to Schmitt and Vogel.
\begin{lemma}\label{Schmitt}{\rm [\cite{SV}, p.~249]} 
Let $P$ be a finite subset of elements of $R$. Let $P_0,\dots, P_r$ be subsets of $P$ such that
\begin{list}{}{}
\item[(i)] $\bigcup_{l=0}^rP_l=P$;
\item[(ii)] $P_0$ has exactly one element;
\item[(iii)] if $p$ and $p''$ are different elements of $P_l$ $(0<l\leq r)$ there is an integer $l'$ with $0\leq l'<l$ and an element $p'\in P_{l'}$ such that $(pp'')^m\in(p')$ for some positive integer $m$.
\end{list}
\noindent
We set $q_l=\sum_{p\in P_l}p^{e(p)}$, where $e(p)\geq1$ are arbitrary integers. We will write $(P)$ for the ideal of $R$ generated by the elements of $P$.  Then we get
$$\sqrt{(P)}=\sqrt{(q_0,\dots,q_r)}.$$
\end{lemma}
This result, together with its refinements and generalizations established in \cite{B2}, is especially useful for ideals generated by monomials. An interesting class of s.t.c.i monomial ideals was introduced by Lyubeznik \cite{L} and studied in \cite{B1}. 
\newline
Big classes of s.t.c.i. ideals generated by binomials have been characterized among the toric ideals (see, e.g., \cite{BL}, \cite{BMT}, \cite{K} and \cite{Br}, \cite{El}, \cite{E}, \cite{P} for toric curves), and some cases are also known among the determinantal ideals  of two-row matrices. One of these was treated by Robbiano and Valla \cite{RV1}, another one by Bardelli and Verdi \cite{BV}.  In this paper we generalize Bardelli and Verdi's result by presenting a class of s.t.c.i. ideals generated by some minors  and some products of entries of certain blockwise defined matrices, which were considered in \cite{BM1}. The corresponding varieties include the fiber cones of codimension two monomial varieties, which were studied by Gim\'enez, Morales and Simis in \cite{G} and  \cite{GMS}, and, furthermore, they belong to the (larger) class of reducible varieties of minimal degree classified by Xamb\'o \cite{X}. These are all defined by monomials and binomials, but they are not all s.t.c.i.. At the end we shall exhibit a counterexample, which is also interesting from another point of view: in positive characteristics its arithmetical rank is strictly greater than its cohomological dimension, which seems to be a rare property.  
\section{A class of set-theoretic complete intersections}
Let $r$ be a positive integer and consider the two-row matrix
\begin{equation}\label{2}A=\left(B_1\|B_2\|\dots\|B_r\right),\end{equation}
\noindent
where, for all $i=1,\dots, r$, $B_i$ is the $2\times c_i$-matrix
\begin{equation}\label{2'}B_i=\left(\begin{array}{ccccc}
X^i_1&X^i_2&\dots&X^i_{c_i-1}&X^i_{c_i}\\
X^i_2&X^i_3&\dots&X^i_{c_i}&X^i_{c_i+1}
\end{array}
\right).
\end{equation}
\noindent
Here \underline{$X$}$=\{X^i_j\}$ is a set of $N$ indeterminates over $K$, and $X^i_j\ne X^h_k$ for $(i,j)\ne(h,k)$, with one only possible exception: for every index $i$, $1\leq i<r$, there is at most one index $i'>i$ such that $X^i_{c_i+1}=X^{i'}_1$. This kind of matrix was introduced by Gim\'enez \cite{G}, and also considered in \cite{BM1}, where $A$ was called a {\it barred matrix} and the $B_i$'s were called the {\it big blocks} of $A$. 
\begin{example}\label{Example1}{\rm  An example of matrix of type (\ref{2}) in the indeterminates $X_1,\dots, X_{11}$ is the following:
$$A=\left(\begin{array}{ccccccccccccc}
X_1&\vline\, \vline&X_3&X_4&\vline\, \vline&X_5&X_6&X_7&\vline\, \vline&X_9&X_{10}\\
X_2&\vline\, \vline&X_4&X_5&\vline\, \vline&X_6&X_7&X_8&\vline\, \vline&X_{10}&X_{11}
\end{array}\right)
$$}
\end{example}
Fix one index $i$, $1\leq i\leq r$. If $c_i>1$, for all indices $j,j'$, $1\leq j<j'\leq c_i$ let $M^i_{jj'}$ be the minor of $B_i$ formed by the $j$-th and the $j'$-th column, i.e., 
$$M^i_{jj'}=\left|\begin{array}{cc}
X^i_j&X^i_{j'}\\
X^i_{j+1}&X^i_{j'+1}
\end{array}
\right|
=X^i_jX^i_{j'+1}-X^i_{j'}X^i_{j+1}.$$
\noindent Let $I_i\in K[X^i_1,\dots, X^i_{c_i+1}]$ be the ideal generated by all minors $M^i_{jj'}$, and consider 
\begin{equation}\label{3}
\!\!\!\!\!\!\!\! F^i_j=\displaystyle\sum_{k=0}^{j}(-1)^k{{j}\choose k}(X^i_{j+2})^{j-k}X^i_{k+1}(X^i_{j+1})^k,\quad\mbox{ for } j=1,\dots, c_i-1,
\end{equation}
\noindent
The first part of the next theorem can be found in \cite{EG}, pp.~118ff, the second part is due to Bardelli and Verdi \cite{BV} (see  \cite{RV2}, Section 2). In order to simplify the notation, we shall omit the index $i$ in the claim.
\begin{theorem}\label{theorem1} The ideal $I$ is prime of height $c-1$. It is a s.t.c.i. on $F_1, \dots, F_{c-1}$. 
\end{theorem}
The ideal $(I_i)$  thus defines an irreducible curve in ${\bf P}_K^{c_i}$, which is known as a {\it rational normal curve}, and is a special kind of {\it (rational normal) scroll}: we refer to Harris \cite{H} for an introductory treatment of this notion. A matrix of type (\ref{2'}) will be called a {\it scroll matrix}. As a consequence of Theorem \ref{theorem1} and Hilbert Nullstellensatz we also have that
\begin{equation}\label{4}\sqrt{(F^i_1,\dots,F^i_{c_i-1})}=I_i.\end{equation}
\noindent
 Following \cite{BM1}, we can associate with matrix $A$ an ideal $J$ of $R=K[$\underline{$X$}$]$. This ideal $J$ is generated by the union of
\begin{list}{}{}
\item[(i)] the set of all minors $M^i_{jj'}$  with $i=1,\dots, r$, and $1\leq j<j'\leq c_i$ (we set $M^i_{jj'}$ equal to the empty set whenever $c_i=1$);
\item[(ii)] the set of all products $X^i_jX^{i'}_{j'}$, with $1\leq i<i'\leq r$, $1\leq j\leq c_i$, $2\leq j'\leq c_{i'}+1$. These are the products of one entry of the upper row of $B_i$ and one entry of the lower row of one of the blocks $B_{i'}$ following on the right.  
\end{list}
Note that 
\begin{equation}\label{5}I_i\subset J\qquad\qquad\mbox{for all }i=1,\dots, r\end{equation}
\noindent
\begin{example}\label{Example2}{\rm The ideal $J$ associated with the matrix $A$ given in Example \ref{Example1} is generated by the following 28 elements
\begin{list}{}{}
\item[(i)] $M^2_{12}=X_3X_5-X_4^2$,\vskip.1truecm\noindent\hphantom{(i)}$M^3_{12}=X_5X_7-X_6^2$,  $M^3_{23}=X_6X_8-X_7^2$,  $M^3_{13}=X_5X_8-X_6X_7$,
\vskip.1truecm\noindent\hphantom{(i)}$M^4_{12}=X_{9}X_{11}-X_{10}^2$;
\item[(ii)]$X_1X_4$, $X_1X_5$, $X_1X_6$, $X_1X_7$, $X_1X_8$, $X_1X_{10}$, $X_1X_{11}$,
\vskip.1truecm\noindent\hphantom{(ii)}$X_3X_6$, $X_3X_7$, $X_3X_8$, $X_3X_{10}$, $X_3X_{11}$,
\vskip.1truecm\noindent\hphantom{(ii)}$X_4X_6$, $X_4X_7$, $X_4X_8$, $X_4X_{10}$, $X_4X_{11}$,
\vskip.1truecm\noindent\hphantom{(ii)}$X_5X_{10}$, $X_5X_{11}$, $X_6X_{10}$, $X_6X_{11}$, $X_7X_{10}$, $X_7X_{11}$.
\end{list}}
\end{example}
We introduce one more piece of notation. We set
\begin{equation}\label{6}G_k=\sum_{i=1}^{r-k}X^i_1X^{i+k}_{c_{i+k}+1},\qquad\mbox{for all }k=1,\dots, r-1.\end{equation}
\noindent
In other words, $G_k$ is the sum of products of the entry lying in the left upper corner of a block $B_i$ and the entry lying in the right lower corner of the block $B_{i+k}$. In particular every summand of $G_k$ is an element of the set defined in (ii), therefore
\begin{equation}\label{7} G_k\in J\qquad\mbox{for all }k=1,\dots, r-1.\end{equation}
\noindent
We are now ready to state our main result, which generalizes Theorem \ref{theorem1}.
\begin{theorem}\label{theorem2} The ideal $J$ is reduced of pure height $\sum_{i=1}^r c_i-1$. It is a s.t.c.i. on 
$$F^1_1,\dots, F^1_{c_1-1},\dots, F^r_1,\dots,F^r_{c_r-1},G_1,\dots, G_{r-1}.$$
\end{theorem}
\demo The first part of the claim is Corollary 1.3 in \cite{BM1}. We prove the second part. First of all we remark that the number of polynomials $F^i_j, G_k$ is
$$\sum_{i=1}^r (c_i-1)+r-1 = \sum_{i=1}^r c_i -r+r-1=\sum_{i=1}^r c_i-1.$$
\noindent 
Next we have to show that, in $K^n$, the variety $V$ defined by $J$ coincides with the variety $V'$ defined by the polynomials $F^i_j, G_k$. According to (\ref{4}), (\ref{5}) and (\ref{7}), all these polynomials belong to $J$, so that we certainly have $V\subset V'$. For the opposite inclusion, consider a point ${\bf x}\in K^n$ such that 
\begin{eqnarray} F^i_j({\bf x})&=&0\qquad\mbox{for all }i=1,\dots, r, j=1,\dots, c_{i}-1,\label{8}\\
G_k({\bf x})&=&0\qquad\mbox{for all }k=1,\dots, r-1.\label{9}
\end{eqnarray}
\noindent
Our claim is that all polynomials in the sets defined in (i) and (ii) vanish at ${\bf x}$.  By virtue of Theorem \ref{theorem1}, (\ref{8}) implies that, for all $i=1,\dots, r$ and $1\leq j<j'\leq c_i$,
\begin{equation}\label{10} M^i_{jj'}({\bf x})=0.\end{equation}
\noindent
 Thus there remains to prove that, for all $i,j, i', j'$ with $1\leq i<i'\leq r, 1\leq j\leq c_i, 2\leq j'\leq c_{i'}+1$,
\begin{equation}\label{11}x^i_jx^{i'}_{j'}=0,\end{equation}
\noindent
where $x^i_j$ is the monomial $X^i_j$ evaluated at ${\bf x}$.
Note that, according to (\ref{6}), 
$$G_{r-1}=X^1_1X^r_{c_r+1}.$$
\noindent
Let $1\leq k<r-1$, and consider the product of two arbitrary summands of $G_k$:
$$\Pi=X^i_1X^{i+k}_{c_{i+k}+1}\cdot X^{i'}_1X^{i'+k}_{c_{i'+k}+1},$$
\noindent
where $1\leq i<i'\leq r-k$. Let $k'$ be such that $i'+k=i+k'$. Then $k'>k$, and $X^{i}_1X^{i'+k}_{c_{i'+k}+1}=X^{i}_1X^{i+k'}_{c_{i+k'}+1}$ is a summand of $G_{k'}$ which divides $\Pi$. Let for all $k=0,\dots, r-2$, $P_k$ be the set of all summands of $G_{r-1-k}$: then Lemma \ref{Schmitt} applies. Thus (\ref{9}), together with Hilbert Nullstellensatz, implies that \begin{equation}\label{12}x^i_1x^{i+k}_{c_{i+k}+1}=0, \qquad\mbox{for all }i=1,\dots, r-1, k=1,\dots, r-i.\end{equation}
\noindent
Now suppose for a contradiction that (\ref{11}) is false, i.e., that
 for some $i,j,j',k$ with $1\leq i\leq r-1, 1\leq k\leq r-i, 1\leq j\leq c_i, 2\leq j'\leq c_{i+k}+1$,  \begin{equation}\label{13}x^i_jx^{i+k}_{j'}\neq0.\end{equation}
\noindent
For two fixed indices $i$ and $k$, let $j'-j$ be maximal with respect to (\ref{13}). Then, by (\ref{12}), necessarily $j'-j\ne c_{i+k}$, which means $j'-j<c_{i+k}$, i.e., $1<j\leq c_i$ or 
$2\leq j'<c_{i+k}+1$. First suppose that $1<j\leq c_i$. Then, by the maximality condition, we have
\begin{equation}\label{14} x^i_{j-1}x^{i+k}_{j'}=0, \end{equation}
\noindent which, in view of (\ref{13}), implies that
\begin{equation}\label{15} x^i_{j-1}=0. \end{equation}
\noindent
On the other hand, by (\ref{10}),
\begin{equation}\label{16} 
M^i_{j-1j}({\bf x})=\left|\begin{array}{cc}
x^i_{j-1}&x^i_j\\
x^i_j&x^i_{j+1}
\end{array}
\right|
=x^i_{j-1}x^i_{j+1}-({x^i_j})^2=0.
\end{equation}
\noindent
But (\ref{15}) and (\ref{16}) give $x^i_j=0$, against our assumption (\ref{13}). Now suppose that $2\leq j'<c_{i+k}+1$. In this case 
$$ x^i_jx^{i+k}_{j'+1}=0, $$
\noindent whence, by (\ref{13})
\begin{equation}\label{18} x^{i+k}_{j'+1}=0. \end{equation}
\noindent
Now, by (\ref{10}),
$$
M^{i+k}_{j'-1j'}({\bf x})=\left|\begin{array}{cc}
x^{i+k}_{j'-1}&x^{i+k}_{j'}\\
x^{i+k}_{j'}&x^{i+k}_{j'+1}
\end{array}
\right|
=x^{i+k}_{j'-1}x^{i+k}_{j'+1}-({x^{i+k}_{j'}})^2=0,
$$
\noindent
which, by (\ref{18}), yields $x^{i+k}_{j'}=0$. This, again, contradicts (\ref{13}). Hence (\ref{11}) is always true. This completes the proof.
\begin{example}\label{example3}{\rm  According to Theorem \ref{theorem2}, the ideal $J$ defined in Example \ref{Example2} has pure height 7 and is a s.t.c.i. on the following polynomials:
\begin{eqnarray*}
F^2_1&=&X_3X_5-X_4^2\\
F^3_1&=&X_5X_7-X_6^2\\
F^3_2&=&X_5X_8^2-2X_6X_7X_8+X_7^3\\
F^4_1&=&X_9X_{11}-X_{10}^2\\
G_3&=&X_1X_{11}\\
G_2&=&X_1X_8+X_3X_{11}\\
G_1&=&X_1X_5+X_3X_8+X_5X_{11}.
\end{eqnarray*}
Computations with  CoCoA \cite{cocoa} show that, for all fields $K$, the least power of $J$ contained in the ideal $(F^2_1,
F^3_1, F^3_2, F^4_1, G_1, G_2, G_3)$ is the 13th. 
}
\end{example}

\section{The fiber cone of a monomial variety of codimension two}
In this section we show that Theorem \ref{theorem2} applies to a relevant class of polynomial ideals. We first need to recall some preliminary notions from commutative algebra. Let $I$ be an ideal of $R$, and let $t$ be an indeterminate over $R$. The graded ring $R[It]=\oplus_{i\in{\bf N}}I^it^i$ is called the {\it Rees algebra} of $I$. If ${\cal M}$ is the ideal of $R$ generated by the set of indeterminates \underline{$X$}, so that $R/{\cal M}\simeq K$, then the quotient ring $F(I)=R[It]/MR[It]\simeq R[It]\otimes K$ is called the {\it fiber cone algebra}  of $I$. Suppose that $H_1,\dots, H_m$ form a minimal generating set of $I$, and introduce  a set of $m$ independent variables over $R$, say \underline{$T$}=$\{T_1,\dots, T_m\}$. Consider the ring homomorphism 
$$\phi:R[\mbox{\underline{$T$}}]\to R[It]$$
\noindent
such that, for all $i=1,\dots, m$,
$$\phi(T_i)=H_it.$$
\noindent 
It is evidently surjective, so that $R[It]\simeq R[\mbox{\underline{$T$}}]/\mbox{Ker}\,\phi$. Tensoring with $K$ yields $F(I)\simeq K[\mbox{\underline{$T$}}]/J$, for a suitable presentation ideal $J\subset K[\mbox{\underline{$T$}}]$.\par\smallskip\noindent
  Assume that $I\subset R=K[X_1,\dots, X_n, Y_1, Y_2]$ is the defining ideal of a {\it a  monomial variety of codimension two}, i.e., of a variety of $K^{n+2}$ admitting a parametrization of the following form:
$$x_1=u_1^{a_1},\ x_2=u_1^{a_2},\dots, \ x_n=u_1^{a_n}, \ y_1=u_1^{b_1}u_2^{b_2}\cdots u_n^{b_n}, \ y_2=u_1^{c_1}u_2^{c_2}\cdots u_n^{c_n},$$
\noindent
where, $a_1,a_2,\dots, a_n$ are positive integers and, for all $i=1,\dots, n$, the exponents $b_i, c_i$ are non negative integers such that $(b_i, c_i)\ne(0,0)$,  and, moreover, $(b_1,b_2,\dots, b_n)\ne(0,\dots, 0)$ and $(c_1,c_2,\dots, c_n)\ne(0,\dots, 0)$. This is an example of affine toric variety. If the above parametrization is homogeneous, it defines a projective variety of ${\bf P}^{n+1}$. 
The presentation ideal $J$ of the fiber cone algebra $F(I)$ is known to be of the type discussed in Section 1 (see \cite{BM1}, Proposition 3.6). An explicit construction of the barred matrix $A$ associated with $J$ can be found in \cite{G} or \cite{GMS}. We present an example which was considered, from a different point of view, in \cite{BM1}, Example 3.7 (b).  
\begin{example}\label{example4} {\rm Consider the projective monomial curve of ${\bf P}^3$ parametrized by
$$x_1=u_1^{534}, \ x_2=u_2^{534}, \ y_1=u_1^{245}u_2^{289}, \ y_2=u_1^{144}u_2^{390}$$
\noindent
Its defining ideal $I\subset R=K[X_1,X_2, Y_1, Y_2]$ is minimally generated by the following 6 binomials:
\begin{eqnarray*}
P_1&=&y_1^{42}-x_1^{19}x_2^{22}y_2\\
P_2&=&y_1^{12}x_2^{3}-x_1^2y_2^{13}\\
P_3&=&y_1^{30}y_2^{12}-x_1^{17}x_2^{25}\\
P_4&=&y_1^{18}y_2^{25}-x_1^{15}x_2^{28}\\
P_5&=&y_1^6y_2^{38}-x_1^{13}x_2^{31}\\
P_6&=&y_2^{51}-y_1^6x_1^{11}x_2^{34},
\end{eqnarray*}
\noindent
and the presentation ideal $J\subset K[T_1,\dots, T_6]$ of $F(I)$ is associated with the barred matrix
$$\left(\begin{array}{cccccc}
T_1&\vline\,\vline&T_3&T_4&\vline\,\vline&T_5\\
T_2&\vline\,\vline&T_4&T_5&\vline\,\vline&T_6
\end{array}
\right).
$$
\noindent
Therefore $J=(T_3T_5-T_4^2, \ T_1T_4,\ T_1T_5,\ T_1T_6,\ T_3T_6,\ T_4T_6,)$, and $J$ is a s.t.c.i. on the following 3 polynomials:
\begin{eqnarray*}F_1&=&T_3T_5-T_4^2\\
G_2&=&T_1T_6\\
G_1&=&T_1T_5+T_3T_6.
\end{eqnarray*}
It can be easily checked that $I^2\subset (F_1, G_1, G_2)$.}
\end{example}
\section{On varieties of minimal degree}
The classifcation of projective varieties of minimal degree (i.e., of degree equal to the codimension plus one, see \cite{EG} for the details) is due to the contributions of various authors. The irreducible ones are the quadric hypersurfaces and the cones over the Veronese surface in ${\bf P}^5$. The reducible case was settled by Xamb\'o, who proved the following
\begin{theorem}\label{theorem3}{\rm [\cite{X}, Section 1]}
Let $J$ be a reduced homogeneous ideal of $R$, defining a reducible variety $V$ of ${\bf P}^n$ of pure dimension $d$ and degree $\delta\leq n-d+1$. Then ${\bf P}^n$ contains linear subspaces $L_1,\dots, L_r$ and there are $d$-dimensional scrolls $V_i\subset L_i$ such that $V=\bigcup_{i=1}^rV_i$ and for each $i=2,\dots, r$, 
$$V_i\cap(V_1\cup\cdots\cup V_{i-1})=L_i\cap(\overline{L_1\cup\cdots\cup L_{i-1}}),$$
\noindent
which is a linear subspace of dimension $d-1$ (the bar denotes the projective closure). Moreover, $\delta=n-d+1$. 
\end{theorem}
The original version of the theorem contained the assumption of connectivity in codimension one, but this can be dropped, as shown in \cite{BM3}. A constructive characterization of the defining ideals of the varieties in Theorem \ref{theorem3} was given in \cite{BM2}. They include {\it all} the ideals studied in \cite{BM1}, i.e., not only the ideals of the type discussed in Section 1, but also the ideals associated to a more general kind of barred matrix,
\begin{eqnarray*}&&\!\!\!\!\!\!\!\!A=\\
&&\!\!\!\!\!\!\!\!\!\!\!\!\left(\begin{array}{ccccccccccccccccc}
B_{11}&\vline& B_{12}&\vline\ \cdots\ \vline& B_{1s_1}&\vline\,\vline&B_{21}&\vline\ \cdots\ \vline& B_{2s_2}&\vline\,\vline\ \cdots\ \vline\,\vline&B_{r1}&\vline\ \cdots\ \vline& B_{rs_r}
\end{array}\right),
\end{eqnarray*}
\noindent
 where every big block $B_i$ may consist of various {\it small blocks} $B_{i1}, \dots,B_{is_i}$, which are scroll matrices and have pairwise distinct sets of entries. We show that the s.t.c.i. property established in Theorem \ref{theorem2} extends to some, but not to all ideals associated with this larger class of barred matrices. We preliminarily remark that Theorem \ref{theorem1} is true also in this more general case (see \cite{BM1}, Corollary 1.3). 
\begin{example}\label{example4'}{\rm Let 
$$A=\left(\begin{array}{ccccc}
X_1&\vline& X_3&\vline\, \vline& X_4\\
X_2&\vline& X_4&\vline\, \vline& X_5
\end{array}\right).$$
\noindent
The associated ideal is 
$$J=(X_1X_4-X_2X_3, \ X_1X_5, \ X_3X_5),$$
\noindent
and height\,$J=2$. We show that $J$ is a s.t.c.i. on the following two polynomials
\begin{eqnarray*} P_1&=&(X_1X_4-X_2X_3)X_2+X_1X_5\\
P_2&=&(X_1X_4-X_2X_3)X_4+X_3X_5,
\end{eqnarray*}
\noindent
i.e., that
\begin{equation}\label{rad}J=\sqrt{(P_1,P_2)}.\end{equation}
\noindent
It holds:
\begin{eqnarray*}
(X_1X_5)^2&=&X_1(X_5-X_2X_4)P_1+X_1X_2^2P_2\\
(X_3X_5)^2&=&X_3(X_5+X_2X_4)P_2-X_3X_4^2P_1,
\end{eqnarray*}
\noindent
so that 
\begin{equation}\label{rad1}X_1X_5,\, X_3X_5\in \sqrt{(P_1,P_2)},\end{equation}
whence, by definition of $P_1$ and $P_2$,
$$(X_1X_4-X_2X_3)X_2, \ (X_1X_4-X_2X_3)X_4\in \sqrt{(P_1,P_2)}.$$
\noindent
It follows that 
$(X_1X_4-X_2X_3)^2\in\sqrt{(P_1,P_2)}$, i.e.,
\begin{equation}\label{rad2}X_1X_4-X_2X_3\in\sqrt{(P_1,P_2)}.\end{equation}
\noindent
Relations (\ref{rad1}) and (\ref{rad2}) imply (\ref{rad}).}
\end{example}
In the next example, the s.t.c.i. property fails to be true.
\begin{example}\label{example5}{\rm Let 
$$A=\left(\begin{array}{ccccccc}
X_1&\vline& X_3&\vline& X_5&\vline\,\vline&X_6\\
X_2&\vline& X_4&\vline& X_6&\vline\,\vline&X_7
\end{array}\right).$$
\noindent
The associated ideal is 
$$J=(X_1X_4-X_2X_3,\ X_3X_6-X_4X_5, \ X_1X_6-X_2X_5, \ X_1X_7, \ X_3X_7, \ X_5X_7),$$
\noindent
 and height\,$J=3$. 
}
\end{example} Let ara\,$J$ denote the arithmetical rank of $J$. 
We show that   ara\,$J>\,$height\,$J$.  We shall use the following criterion, which is based on \'etale cohomology and is due to Newstead \cite{N}.
\begin{lemma}\label{Newstead}{\rm [\cite{BS}, Lemma 3$^\prime$]} Let
$W\subset\tilde W$ be affine varieties. Let $d=\dim\tilde
W\setminus W$. If there are $s$ equations $F_1,\dots, F_s$ such
that $W=\tilde W\cap V(F_1,\dots,F_s)$, then 
$$\het^{d+i}(\tilde W\setminus W,{\bz}/r{\bz})=0\quad\mbox{ for all
}i\geq s$$ and for all $r\in{\bz}$ which are prime to {\rm char}\,$K$.
\end{lemma}
We refer to \cite{Mi} for the basic notions on \'etale cohomology. 
Let $p$ be a prime such that $p\ne$\,char\,$K$. Let $V=V(J)\subset K^7$. In view of Lemma \ref{Newstead}, for our purpose it suffices to show that 
\begin{equation}\label{19} \het^{10}(K^7\setminus V, {\bz}/p{\bz})\ne0.
\end{equation}
\noindent
 By Poincar\'e Duality (see \cite{Mi}, Theorem 14.7, p.~83) we have
\begin{equation}\label{20} {\rm Hom}_{\scriptstyle{\bz}/p{\bz}} (\het^{10}(K^7\setminus V, {\bz}/p{\bz}), {\bz}/p{\bz})\simeq  \hc^{4}(K^7\setminus V, {\bz}/p{\bz}),
\end{equation}
\noindent
where $\hc$ denotes \'etale cohomology with compact support. 
For the sake of simplicity, we shall omit the coefficient group ${\bz}/p{\bz}$ henceforth. Let $W$ be the subvariety of $K^7$ defined by 
$$X_1X_4-X_2X_3=0,\ X_3X_6-X_4X_5=0,\ X_1X_6-X_2X_5=0, \ X_7=0.$$
\noindent
Then $W\subset V$, and 
$$V\setminus W=\{(x_1,\dots, x_7)\in K^7\vert x_1=x_3=x_5=0, \ x_7\ne0\}\simeq K^3\times (K\setminus\{0\}).$$
\noindent
It is well-known that
\begin{equation}\label{21}\hc^i(K^t)\simeq\left\{\begin{array}{cl} {\bz}/p{\bz}&\mbox{if }i=2t\\
0&\mbox{else, }
\end{array}\right.
\end{equation}
\noindent
and 
\begin{equation}\label{22}\hc^i(K^t\setminus\{0\})\simeq\left\{\begin{array}{cl} {\bz}/p{\bz}&\mbox{if }i=1,2t\\
0&\mbox{else. }
\end{array}\right.
\end{equation}
\noindent
Moreover, by the K\"unneth formula (\cite{Mi}, Theorem 22.1), 
$$\hc^i(V\setminus W)\simeq\displaystyle\bigoplus_{h+k=i}\hc^h(K^3)\otimes\hc^k(K\setminus\{0\}),$$
\noindent
so that, by (\ref{21}),
\begin{equation}\label{23}\hc^3(V\setminus W)=\hc^4(V\setminus W)=0.\end{equation}
We have a long exact sequence of \'etale cohomology with compact support:
$$\cdots\rightarrow\hc^3(V\setminus W)\rightarrow \hc^4(K^7\setminus V)\rightarrow \hc^4(K^7\setminus W)\rightarrow 
\hc^4(V\setminus W)\rightarrow\cdots.$$
\noindent
By (\ref{23}) it follows that
\begin{equation}\label{24} \hc^4(K^7\setminus V)\simeq\hc^4(K^7\setminus W)
\end{equation}
\noindent
In view of (\ref{20}) and (\ref{24}) our claim (\ref{19}) will follow once we have proven that 
\begin{equation}\label{25} \hc^4(K^7\setminus W)\ne0.\end{equation}
\noindent
 This is what we are going to prove next. Note that $W$ is the variety of $K^6$ defined by the vanishing of the 2-minors of a generic $2\times 3$ matrix of indeterminates. Thus $\bar W=W\setminus\{0\}$ is the set of non zero 2$\times$3 matrices with proportional rows. The set of such matrices where the first row is zero is a closed subset of $\bar W$ which can be identified with $Z=K^3$, and its complementary set is $\bar W\setminus Z\simeq K^3\setminus\{0\}\times K$.  We thus have a
long exact sequence of \'etale cohomology with compact support:
\begin{equation}\label{27a}\cdots\rightarrow\hc^2(Z)\rightarrow \hc^3(\bar W\setminus Z)\rightarrow \hc^3(\bar W)\rightarrow \hc^3(Z)\rightarrow\cdots,\end{equation}
\noindent
where, according to (\ref{21}), $\hc^2(Z)=\hc^3(Z)=0$, whereas, by  the K\"unneth formula and (\ref{22}), 
$$\hc^3(\bar W\setminus Z)\simeq \hc^1(K^3\setminus\{0\})\otimes\hc^2(K)\simeq \bz/p\bz.$$
\noindent
It follows that (\ref{27a}) gives rise to an isomorphism
\begin{equation}\label{26}\hc^3(\bar W)\simeq{\bz}/p{\bz}.\end{equation}
\noindent
On the other hand there is also the following long exact sequence of \'etale cohomology with compact support:
\begin{equation}\label{27}\cdots\rightarrow\hc^3(K^7\setminus \{0\})\rightarrow \hc^3(\bar W)\rightarrow \hc^4(K^7\setminus W)\rightarrow 
\cdots,\end{equation}
\noindent
where 
\begin{equation}\label{28} \hc^3(K^7\setminus\{0\})=0.\end{equation}
\noindent
by (\ref{22}). Hence, in view of (\ref{28}) and (\ref{26}), in (\ref{27}) the left term is zero, and the middle term is non zero. It follows that the right term is non zero, i.e., claim (\ref{25}) holds. This proves that ara\,$J>3$, as desired.
It can be easily checked that the variety $V$ is defined by the following 5 equations:
$$X_1X_4-X_2X_3=0,\ X_3X_6-X_4X_5=0,$$
$$X_1X_6-X_2X_5+X_3X_7=0, \ X_1X_7=0, \ X_5X_7=0.$$ 
\noindent
Hence ara\,$J\leq5$. We conjecture that ara\,$J=5$. 
\par\medskip\noindent
\'Etale cohomology is not the only possible tool for finding a lower bound for the arithmetical rank. In general we have
\begin{equation}\label{29}{\rm cd}\,J\leq{\rm ara}\,J,\end{equation}
\noindent where 
$${\rm cd}\,J=\max\{i\vert H^i_J(R)\ne0\},$$
\noindent 
is called the {\it cohomological dimension} of $J$. Here $H^i_J$ denotes the $i$-th local cohomology group with respect to $J$;  we refer to Huneke \cite{Hu} for an extensive exposition of this subject. Inequality (\ref{29}) can be strict, but not many examples  of this kind are known so far. One monomial ideal was found by Zhao Yan \cite{Z}, Example 2 following \cite{SV}, p.~250, and \cite{L}, Example 1,  whereas classes of determinantal ideals are described in \cite{BS} and \cite{B0}. In all these cases (\ref{29}) is strict in all but one characteristics. We show that for the ideal $J$ discussed in Example \ref{example5} we always have cd\,$J<5$: if our above conjecture is true, this would provide an example of ideal whose cohomological dimension differs from its arithmetical rank in all characteristics.  According to \cite{Hu}, Theorem 2.2, we have the following long exact sequence of local cohomology:
\begin{equation}\label{30} 
\cdots\rightarrow H^i_{J+(X_7)}(R)\rightarrow H^i_J(R)\rightarrow H^i_J(R_{X_7})\rightarrow\cdots,\end{equation}
\noindent
where, by \cite{Hu}, Proposition 1.10, $H^i_J(R_{X_7})\simeq H^i_{J_{X_7}}(R_{X_7})$. Now
\begin{eqnarray*} J+(X_7)&=&(X_1X_4-X_2X_3,\ X_1X_6-X_2X_5,\ X_3X_6-X_4X_5,\ X_7)\subset R,\\
J_{X_7}&=&(X_1,\ X_3,\ X_5)\subset R_{X_7}.
\end{eqnarray*}
\noindent
Since $J+(X_7)$ and $J_{X_7}$ are generated by sets of at most 4 elements in $R$ and $R_{X_7}$ respectively, inequality (\ref{29}) implies that
$$H^i_{J+(X_7)}(R)=H^i_{J_{X_7}}(R_{X_7})=0,\qquad\mbox{for all }i\geq5,$$
\noindent
so that, in view of (\ref{30}), 
$$H^i_J(R)=0\qquad\mbox{for all }i\geq5.$$
\noindent
Hence cd\,$J<5$, as was to be shown.  
\begin{remark}{\rm According to \cite{EG}, Theorem 4.2, every variety fulfilling the assumption of Theorem \ref{theorem3} is Cohen-Macaulay. By \cite{PS}, Prop. (4.1), this implies that cd\,$J=\,$ht\,$J=3$. Hence we certainly have that cd\,$J<$ara\,$J$ in all positive characteristics. 
}\end{remark}

\end{document}